How many moles of gaseous boron trifluoride, BF3, are contained in a 434.10-mL bulb at 788.0 K if the pressure is 1.220 torr? How many grams of BF3? Which law is this? Pay attention to the units.
Answer #1

Using the ideal gas equation : n = PV / RT

Where P is pressure in torr, T is temperature in K and V is olume in liters

P = 1.22 torr = 1.22 / 760 atm

n = 1.22 / 760 * 0.434 / (0.0821 * 788) = 1.077 * 10-5 moles

According to Avogadro's law, mass of BF3 = moles * molar mass = 1.077 * 10-5 * 67.8 = 7.30 * 10-4 g